Can you please help me. I have been out of school for over 30 years and am trying to help my child. She had this question on a test. How do I show her how to work out the answer.
You have a baseball diamond that is 11 x 11.How far would Susie have the throw the ball to Mary who is on second base. I said 11 ft but my daughter said 22 ft but both were wrong.
Thank you for your help
Paula

Answer by whisper(27)

A baseball diamond that is 11x11 is a square
home base and 2nd base are on opposite corners
The line from home base to 2nd base forms the hypoteneuse of a right triangle with sides from home plate to first base and first base to second base.
Since we have a right triangle we can use the Phythagorean Theorem
Which states that a^2+b^2=c^2 where a and b are the sides a right triangle with hypoteneuse c.
We know:
a=11
b=11
a^2+b^2=c^2
11^2+11^2=c^2
121+121=c^2
242=c^2
Take the square root of both sides:
15.556349=c
Please don't hesitate to ask if you have further questions!
round it to two decimals:
approximately 15.56 feet